# CALT-TH 2023-052

### 15 Dec 2023 v1

@@
[…]
${}^{1}$alter Burke Institute for Theoretical Physics, California Institute of Technology,
@@
${}^{2}$enter for Theoretical Physics, Massachusetts Institute of Technology,
@@
${}^{3}$epartment of Physics, Arizona State University,
@@
Motivated by bit threads, we introduce a new prescription for computing entropy vectors outside the holographic entropy cone. By utilizing cycle flows on directed graphs, we show that the maximum cycle flow associated to any subset of vertices, which corresponds to a subsystem, manifestly obeys purification symmetry. Furthermore, we prove that the maximum cycle flow obeys both subadditivity and strong subadditivity, thereby establishing it as a viable candidate for the entropy associated to the subsystem. Finally, we demonstrate how our model generalizes the entropy vectors obtainable via conventional flows in undirected graphs and hypergraphs.
@@ 1 Introduction
  • Begin with a balanced directed network. [^3] For us, we choose to call a network a graph with edge weights. This is defined precisely later.
@@
  • Partition the boundary vertices $\partial V$ into the set ${\underline{I}}$ and its complement, ${\underline{I}}^{c}\coloneqq\partial V\setminus{\underline{I}}$.
@@
  • The entropy ${\texttt{S}}({\underline{I}})$ is given by the magnitude of the maximum simultaneous cycle flow along all such paths.
Note that this protocol uses cycle flows, which are flows with neither sources nor sinks, on balanced directed networks, which are graphs with directed edges obeying the condition where at every vertex the total weight of ingoing edges equals that of outgoing edges. Additionally, the specific flow pattern is important; we need to know which cycles have flow turned on, not just which edges. As we shall see, this generalization to directed graphs under a cycle flow protocol allows access to a broader set of entropy vectors than the undirected or hyperedge graph protocols. Additionally, the cycle flow protocol retains some of the topological features of the link picture.
@@ 2.1 Quantum Information Theory
Henceforth, SA and SSA will refer to the inequalities (2.8) and (2.9), which can involve the purifier entropy ${\texttt{S}}({\sf{N}}+1)$.
For any choice of ${\underline{I}},{\underline{J}},{\underline{K}}$, the inequalities (2.8) and (2.9) each define a half-space in entropy space. The intersection of all such half-spaces then forms a polyhedral cone in which entropy vectors corresponding to realizable quantum states live. We will refer to this cone as the SSA entropy cone. The quantum entropy cone (QEC) is defined to be the topological closure of the set of all the entropy vectors realizable by quantum states, and was shown in [28] to be a convex cone. Thus, because all such entropy vectors must obey SA and SSA, the QEC lives inside the SSA entropy cone. It is known that for ${\sf{N}}=3$ the QEC actually coincides with the SSA entropy cone. However, for ${\sf{N}}\geq 4$, besides the existence of infinitely many constrained inequalities (inequalities that only apply on subspaces of codimension higher than zero) [29, 30], very little is known about the QEC. [^5] In particular, it remains an open question whether the QEC is even polyhedral for higher parties.
@@ 2.2 Graph Theory
We begin by introducing notation and establishing conventions for relevant graph theory concepts. Consider first a discrete directed graph ${\mathcal{G}}\coloneqq(V,E)$, where $V$ is the collection of vertices and $E$ is the collection of directed edges. [^6] Note that undirected graphs are a special case of directed graphs where between every two vertices $v_{1},v_{2}\in V$ there exists a directed edge from $v_{1}$ to $v_{2}$ and another directed edge from $v_{2}$ to $v_{1}$. Thus, it suffices for us to study directed graphs only. Associated to every directed edge $e\in E$ is a capacity function $c:\,E\to\mathbb{R}_{\geq 0}$, which maps every edge to a non-negative real number known as the edge’s capacity or weight. We can then define a network ${\mathcal{N}}\coloneqq(V,E,c)$ to be a graph $(V,E)$ with a capacity function $c$.
For every directed edge $e$, it is useful to define $s(e)\in V$ to be the source vertex from which the edge originates from, and $t(e)\in V$ to be the target vertex towards which the edge is directed. Using the capacity function, we will now introduce the following notion of cycle flows, which will be useful for our purposes.
@@
Unlike typical flows in network theory involving sources and sinks, which have arisen in quantum information literature as bit threads [11], simple cycle flows have not been connected to any notion of entropy. However, we will introduce a proposal in the next section that relates cycle flows to entanglement entropy beyond holographic states. It would be interesting to see if there are any relations between the cycle flows we are considering and the more well-understood bit threads (see Footnote 13). [^7] For an exploration of the more conventional flows involving sources and sinks, and how they pertain to the study of the holographic entropy cone in the context of bit threads, we refer the reader to [11] and in particular Section 6 of [13].
@@
[…]
@@
In the following section, we will identify the maximum flux of a multi-cycle flow through some boundary vertices ${\underline{I}}$ of a balanced network (defined in 7) as the entropy, hence the suggestive notation of denoting it as ${\texttt{S}}({\underline{I}})$. Furthermore, for convenience, we will henceforth denote
@@
Figure 1: All the edges implicitly are weight 1. We illustrate two multi-cycle flows on the graph. In both cases, the multi-cycle flow consists of two simple cycles, which are colored red and blue in the figure. Notice that the two multi-cycle flows give rise to different fluxes through the subsets of vertices.
@@
Given a partition of the boundary vertices $\partial V=\{{\underline{I}}_{1},\ldots,{\underline{I}}_{n}\}$, a multi-cycle flow ${\mathfrak{C}}$ is minimal with respect to $\{{\underline{I}}_{1},\ldots,{\underline{I}}_{n}\}$ if all the simple cycle flows comprising ${\mathfrak{C}}$ are minimal with respect to $\{{\underline{I}}_{1},\ldots,{\underline{I}}_{n}\}$.
We conclude by proving a result regarding minimal cycles that will later be useful. Henceforth, we will only focus on a class of networks known as balanced networks, which are defined as follows.
@@
Recall now the notion of a residual network given a network ${\mathcal{N}}=(V,E,c)$ and a multi-cycle flow =${\mathfrak{C}}=\{(E_{i};{\lambda}_{i})\}_{i=1,\ldots,m}$. Utilizing the definition of $u_{i}$ given in 3, the residual graph is given by
@@
[…]
@@
[…]
@@
To prove the second part of the claim, consider some multi-cycle flow ${\mathfrak{C}}^{\prime}$ in ${\mathcal{N}}$. It is obvious that the capacity of every edge in ${\mathcal{N}}_{1}\coloneqq{\text{Res}}({\mathcal{N}},{\mathfrak{C}}^{\prime})$ is less than or equal to that in ${\mathcal{N}}$, with the capacity of at least one edge in ${\mathcal{N}}_{1}$ being strictly less than that in ${\mathcal{N}}$. If $c(e)>0$ for some $e\in E$ in ${\mathcal{N}}_{1}$, we can construct a simple cycle ${\mathcal{C}}_{1}$ containing $e$ by first constructing an edge cycle $E_{{\mathcal{C}}_{1}}$ explicitly starting from $e$; this is always possible because ${\mathcal{N}}_{1}$ is balanced. The magnitude of ${\mathcal{C}}_{1}$ is then given to be $\text{min}_{e\in E_{{\mathcal{C}}_{1}}}c(e)$. Notice this means the residual graph ${\mathcal{N}}_{2}\coloneqq{\text{Res}}({\mathcal{N}}_{1},{\mathcal{C}}_{1})$ has at least one fewer edge than ${\mathcal{N}}_{1}$. We can then repeat this process with ${\mathcal{N}}_{1}$ and ${\mathcal{N}}$ and ${\mathcal{N}}_{2}$ in place of ${\mathcal{N}}_{1}$ and obtain a simple cycle flow ${\mathcal{C}}_{2}$. Repeating this until the resulting residual graph has no edges, i.e. $c(e)=0$ for all edges, it then follows ${\mathfrak{C}}\coloneqq{\mathfrak{C}}^{\prime}\cup{\mathcal{C}}_{1}\cup\cdots\cup{\mathcal{C}}_{n}$ is a multi-cycle flow where the residual graph ${\text{Res}}({\mathcal{N}},{\mathfrak{C}})$ has $c(e)=0$ for every edge. ∎
We next prove that cycle flows on a balanced network have a crucial property known as nesting. [^8] This claim is analogous to the result from [11], where they prove the existence of nested flows for conventional flows involving sources and sinks. It would be interesting to further explore connections between cycle flows and conventional flows, especially in the context of multi-commodity flows. As we will see in the next section, this property will be repeatedly used to prove that (2.18) obeys both SA and SSA, thereby motivating us to identify it with the entropy.
@@
Given a balanced network ${\mathcal{N}}=(V,E,c)$ with boundary vertices $\partial V\subseteq V$, and given two disjoint subsets of boundary vertices ${\underline{I}},{\underline{J}}$ where ${\underline{I}}\cap{\underline{J}}=\emptyset$, let ${\underline{K}}\coloneqq\partial V\setminus{\underline{I}}{\underline{J}}$ denote the rest of the boundary vertices. Then there exists a minimal multi-cycle flow ${\mathfrak{C}}({\underline{I}},{\underline{J}})$, with respect to the boundary vertex partition $\partial V=\{{\underline{I}},{\underline{J}},{\underline{K}}\}$ whose magnitude through ${\underline{I}}$ and ${\underline{I}}{\underline{J}}$ is simultaneously maximized.
@@
The proof proceeds by identifying sets of minimal cycles which can be simultaneously combined to construct a multi-cycle flow with the desired properties. We start from a multi-cycle flow ${\mathfrak{C}}_{{\underline{I}}{\underline{J}}{\underline{K}}}$ that maximizes the flux through ${\underline{I}}$ conditioned on only using minimal cycle flows that each include at least one boundary vertex in all ${\underline{I}},{\underline{J}}$, and ${\underline{K}}$. Note that this means ${\mathfrak{C}}_{{\underline{I}}{\underline{J}}{\underline{K}}}$ will already automatically include all contributions to the flux through ${\underline{I}}{\underline{J}}$ which involve at least one vertex in ${\underline{I}}$. By construction, the residual network ${\mathcal{N}}_{1}\coloneqq{\text{Res}}({\mathcal{N}},{\mathfrak{C}}_{{\underline{I}}{\underline{J}}{\underline{K}}})$ no longer contains any minimal cycles involving vertices in all ${\underline{I}},{\underline{J}}$, and ${\underline{K}}$.
@@
Next, on ${\mathcal{N}}_{1}$, let ${\mathfrak{C}}_{{\underline{I}}{\underline{K}}}$ be the multi-cycle flow that maximizes the flux through ${\underline{I}}$ and consists only of minimal cycle flows that each include at least one boundary vertex in both ${\underline{I}}$ and ${\underline{K}}$. By construction, this means that the residual network ${\mathcal{N}}_{2}\coloneqq{\text{Res}}({\mathcal{N}}_{1},{\mathfrak{C}}_{{\underline{I}}{\underline{K}}})$ does not have any minimal cycles simultaneously connecting vertices in both ${\underline{I}}$ and ${\underline{K}}$. In particular, this means any vertex in ${\underline{J}}$ that is part of a minimal cycle with a vertex in ${\underline{K}}$ cannot be connected to any vertices in ${\underline{I}}$, and similarly any vertex in ${\underline{J}}$ that is part of a minimal cycle involving vertices in ${\underline{I}}$ cannot be connected to any vertex in ${\underline{K}}$. It follows that we can simultaneously construct multi-cycle flows ${\mathfrak{C}}_{{\underline{I}}{\underline{J}}}$ and ${\mathfrak{C}}_{{\underline{J}}{\underline{K}}}$ on ${\mathcal{N}}_{2}$ such that ${\mathfrak{C}}_{{\underline{I}}{\underline{J}}}$ only involve minimal cycles that maximize the flux between ${\underline{I}}$ and ${\underline{J}}$, and ${\mathfrak{C}}_{{\underline{J}}{\underline{K}}}$ only involve minimal cycle flows that maximize the flux between ${\underline{J}}$ and ${\underline{K}}$.
Because the networks ${\mathcal{N}}$, ${\mathcal{N}}_{1}$, and ${\mathcal{N}}_{2}$ are all balanced, there are no longer any directed edges left in ${\mathcal{N}}_{3}\coloneqq{\text{Res}}({\mathcal{N}}_{2},{\mathfrak{C}}_{IJ}\cup{\mathfrak{C}}_{JK})$ connecting ${\underline{I}}$, ${\underline{J}}$, or ${\underline{K}}$. The remaining minimal cycles in ${\mathcal{N}}_{3}$ do not contribute to any fluxes through ${\underline{I}}$, ${\underline{J}}$, or ${\underline{K}}$. Hence, it is clear that ${\mathfrak{C}}_{{\underline{I}}{\underline{J}}{\underline{K}}}\cup{\mathfrak{C}}_{{\underline{I}}{\underline{K}}}\cup{\mathfrak{C}}_{{\underline{I}}{\underline{J}}}$ maximizes the flux through ${\underline{I}}$, and ${\mathfrak{C}}_{{\underline{I}}{\underline{J}}{\underline{K}}}\cup{\mathfrak{C}}_{{\underline{I}}{\underline{K}}}\cup{\mathfrak{C}}_{{\underline{J}}{\underline{K}}}$ maximizes the flux through ${\underline{I}}{\underline{J}}$. By our argument in the previous paragraph, since we can simultaneously have ${\mathfrak{C}}_{{\underline{I}}{\underline{J}}}$ and ${\mathfrak{C}}_{{\underline{J}}{\underline{K}}}$, this means the multi-cycle flow ${\mathfrak{C}}_{{\underline{I}}{\underline{J}}{\underline{K}}}\cup{\mathfrak{C}}_{{\underline{I}}{\underline{K}}}\cup{\mathfrak{C}}_{{\underline{I}}{\underline{J}}}\cup{\mathfrak{C}}_{{\underline{J}}{\underline{K}}}$ simultaneously maximizes the flux through both ${\underline{I}}$ and ${\underline{I}}{\underline{J}}$. ∎
Before continuing, we remark that it was crucial in the proof above that we are constructing the multi-cycle flows using minimal cycle flows and not just simple cycle flows. To see why, consider the two simple cycle flows in Figure 2 with ${\underline{I}}=\{A_{1},A_{2}\}$, ${\underline{J}}=\{B\}$, and ${\underline{K}}=\{C\}$. In both cases, the red simple cycle flow maximizes the flux through ${\underline{I}}$ while utilizing at least one vertex in ${\underline{I}},{\underline{J}}$, and ${\underline{K}}$. However, the right figure, which is not a minimal cycle flow, prevents the addition of the minimal cycle involving just the vertices $A_{2}$ and $B$ when constructing ${\mathfrak{C}}_{IJ}$. Therefore, the inclusion of the non-minimal simple cycle depicted in the right network in the construction given in the proof of Theorem 1 will not lead to a multi-cycle flow that maximizes the flux through ${\underline{I}}=\{A_{1},A_{2}\}$.
We conclude this section by noting that the above theorem crucially relies on the fact the network has the balanced property. For instance, consider the unbalanced network below, where every edge has capacity 1. It is easy to see that the maximal flux through $ABC$ is 1, and the maximal flux through $C$ is 2. However, there is only one cycle flow that maximizes the flux through $ABC$, which is shown in red in the figure below, and it does not maximize the flux through $C$. Hence, nesting fails for this unbalanced network.
$O$$C$$B$$A$

### 12 Oct 2024 v3

@@
Temple He1, Sergio Hernández-Cuenca2, Cynthia Keeler3
1Walter Burke Institute for Theoretical Physics, California Institute of Technology,
@@
2Center for Theoretical Physics, Massachusetts Institute of Technology,
@@
3Department of Physics, Arizona State University,
@@
Motivated by bit threads, we introduce a new prescription for computing entropy vectors outside the holographic entropy cone. By utilizing cycle flows on directed graphs, we show that the maximum cycle flow associated to any subset of vertices, which corresponds to a subsystem, manifestly obeys purification symmetry. Furthermore, by restricting ourselves to a subclass of directed graphs, we prove that the maximum cycle flow obeys both subadditivity and strong subadditivity, thereby establishing it as a viable candidate for the entropy associated to the subsystem. Finally, we demonstrate how our model generalizes the entropy vectors obtainable via conventional flows in undirected graphs, as well as conjecture that our model similarly generalizes the entropy vectors arising from hypergraphs.
@@ 1 Introduction
  • Begin with a balanced directed network, which are graphs with directed edges obeying the condition where at every vertex the total weight of ingoing edges equals that of outgoing edges, that also obeys a property known as “nesting” (introduced below). [^3] For us, we choose to call a network a graph with edge weights. This is defined precisely later.
@@
  • Partition the boundary vertices $\partial V$ into the set ${\underline{I}}$ and its complement, ${\underline{I}}^{c}\coloneqq\partial V\setminus{\underline{I}}$. [^4] We underline ${\underline{I}}$ to indicate it may contain the boundary vertex corresponding to the purifier.
@@
  • The entropy ${\texttt{S}}({\underline{I}})$ is given by the magnitude of the maximum simultaneous cycle flow, defined in Section 2.2, along all such paths.
Note that this protocol uses cycle flows, which are flows with neither sources nor sinks, on a subclass of balanced directed networks. Additionally, the specific flow pattern is important; we need to know which cycles have flow turned on, not just which edges. As we shall see, this generalization to directed graphs under a cycle flow protocol allows access to a broader set of entropy vectors than the undirected or hyperedge graph protocols. Additionally, the cycle flow protocol retains some of the topological features of the link picture.
@@ 2.1 Quantum Information Theory
Henceforth, SA and SSA will refer to the inequalities (2.8) and (2.9), which can involve subsystems containing the purifier ${\sf{N}}+1$.
For any choice of ${\underline{I}},{\underline{J}},{\underline{K}}$, the inequalities (2.8) and (2.9) each define a half-space in entropy space. The intersection of all such half-spaces then forms a polyhedral cone in which entropy vectors corresponding to realizable quantum states live. We will refer to this cone as the SSA entropy cone (even though the cone is constrained by both SA and SSA). The quantum entropy cone (QEC) is defined to be the topological closure of the set of all the entropy vectors realizable by quantum states, and was shown in [28] to be a convex cone. Thus, because all such entropy vectors must obey SA and SSA, the QEC lives inside the SSA entropy cone. It is known that for ${\sf{N}}=3$ the QEC actually coincides with the SSA entropy cone. However, for ${\sf{N}}\geq 4$, besides the existence of infinitely many constrained inequalities (inequalities that only apply on subspaces of codimension higher than zero) [29, 30], very little is known about the QEC. [^6] In particular, it remains an open question whether the QEC is even polyhedral for higher parties.
@@ 2.2 Graph Theory
We begin by introducing notation and establishing conventions for relevant graph theory concepts. Consider first a discrete directed graph ${\mathcal{G}}\coloneqq(V,E)$, where $V$ is the collection of vertices and $E$ is the collection of directed edges such that there is at most one directed edge from any vertex $v\in V$ to any other vertex $v^{\prime}\in V$ [^7] Note that undirected graphs are a special case of directed graphs where between every two vertices $v_{1},v_{2}\in V$ there exists a directed edge from $v_{1}$ to $v_{2}$ and another directed edge from $v_{2}$ to $v_{1}$ (see Figure 5). Thus, it suffices for us to study directed graphs only. Associated to every directed edge $e\in E$ is a capacity function $c:\,E\to\mathbb{R}_{\geq 0}$, which maps every edge to a non-negative real number known as the edge’s capacity or weight. We can then define a network ${\mathcal{N}}\coloneqq(V,E,c)$ to be a graph $(V,E)$ with a capacity function $c$.
For every directed edge $e$, it is useful to define $s(e)\in V$ to be the source vertex from which the edge originates from, and $t(e)\in V$ to be the target vertex towards which the edge is directed. Using the capacity function, we will now introduce the following notion of simple cycle flows, which will be useful for our later construction of multi-cycle flows.
@@
Unlike typical flows in network theory involving sources and sinks, which have arisen in quantum information literature as bit threads [11], simple cycle flows have not been connected to any notion of entropy. However, we will introduce a proposal in the next section that relates cycle flows to entanglement entropy beyond holographic states. It would be interesting to see if there are any relations between the cycle flows we are considering and the more well-understood bit threads (see Footnote 17). [^8] For an exploration of the more conventional flows involving sources and sinks, and how they pertain to the study of the holographic entropy cone in the context of bit threads, we refer the reader to [11] and in particular Section 6 of [13].
@@
$f_{\mathcal{C}}({\underline{I}},{\underline{J}})=\begin{cases}\lambda&V({\mathcal{C}})\cap{\underline{I}}\neq\emptyset\quad\text{and}\quad V({\mathcal{C}})\cap{\underline{J}}\neq\emptyset\\ 0&\text{otherwise.}\end{cases}$ (2.13)
@@
In the following section, we will identify the maximum flux of a multi-cycle flow through some boundary vertices ${\underline{I}}$ of a balanced network (defined in 8) as the entropy, hence the suggestive notation of denoting it as ${\texttt{S}}({\underline{I}})$. Furthermore, for convenience, we will henceforth denote
@@
We will also introduce the notion of a subflow in a multi-cycle flow.
Definition 4 (Multi-cycle subflow).
Given a network ${\mathcal{N}}=(V,E,c)$ and a multi-cycle flow :=${\mathfrak{C}}:=\{{\mathcal{C}}_{i}=(E_{i};{\lambda}_{i})\}_{i=1,\ldots,m}$, a multi-cycle =${\mathfrak{C}}^{\prime}=\{{\mathcal{C}}_{i}^{\prime}=(E_{i}^{\prime};{\lambda}_{i}^{\prime})\}_{i=1,\ldots,m}$ is a multi-cycle subflow of ${\mathfrak{C}}$ if $E_{i}^{\prime}=E_{i}$ and ${\lambda}_{i}^{\prime}\leq l_{i}$ for all $i$ [^9] If ${\lambda}_{j}^{\prime}=0$ for some $j$ this just means ${\mathfrak{C}}^{\prime}$ does not contain the simple cycle ${\mathcal{C}}_{j}$.
@@
Figure 1: All the flows implicitly are weight 1. We illustrate two multi-cycle flows on the graph. In both cases, the multi-cycle flow consists of two simple cycles, which are colored red and blue in the figure. Notice that the two multi-cycle flows give rise to different fluxes through the subsets of vertices.
@@
Given a partition of the boundary vertices $\partial V$ into subsets $\{{\underline{I}}_{1},\ldots,{\underline{I}}_{n}\}$, a multi-cycle flow ${\mathfrak{C}}$ is minimal with respect to $\{{\underline{I}}_{1},\ldots,{\underline{I}}_{n}\}$ if all the simple cycle flows comprising ${\mathfrak{C}}$ are minimal with respect to $\{{\underline{I}}_{1},\ldots,{\underline{I}}_{n}\}$.
Finally, we want to begin restricting ourselves to the types of directed networks under consideration. The first restriction we impose is to focus on balanced networks, which are defined as follows.
@@
The reason we want to restrict ourselves to balanced networks is because otherwise, SSA fails if the entropy through some boundary vertices ${\underline{I}}$ is given by the maximum flux through it. For instance, consider the unbalanced network below, where every edge has capacity 1. It is easy to verify that the maximum flux through $C$ is 2, through $AC$ is 1, through $BC$ is 2, and through $ABC$ is 2. But this means ${\texttt{S}}(AC)+{\texttt{S}}(BC)\ngeq{\texttt{S}}(C)+{\texttt{S}}(ABC)$. Hence, SSA fails for this unbalanced network.
$B$$C$$O$$A$
Figure 3: An unbalanced network, where all the edges have capacity 1, that does not satisfy SSA. It is straightforward to verify ${\texttt{S}}(C)=2,{\texttt{S}}(AC)=1,{\texttt{S}}(BC)=2,$ and ${\texttt{S}}(ABC)=2$. Then ${\texttt{S}}(AC)+{\texttt{S}}(BC)=1+2\ngeq{\texttt{S}}(C)+{\texttt{S}}(ABC)=2+2$.
We now introduce the standard notion of a residual network given a network ${\mathcal{N}}=(V,E,c)$, as applied to a multi-cycle flow =${\mathfrak{C}}=\{(E_{i};{\lambda}_{i})\}_{i=1,\ldots,m}$. Utilizing the definition of $u_{i}$ given in 3, the residual network is given by
@@
Given any (not necessarily balanced) network ${\mathcal{N}}=(V,E,c)$ and a multi-cycle flow ${\mathfrak{C}}$, if ${\mathcal{C}}$ is a simple cycle flow in ${\text{Res}}({\mathcal{N}},{\mathfrak{C}})$, then ${\mathfrak{C}}\cup{\mathcal{C}}$ is a multi-cycle flow in the original network ${\mathcal{N}}$.
@@
Given a balanced network ${\mathcal{N}}=(V,E,c)$ and a multi-cycle flow ${\mathfrak{C}}$, the residual network ${\text{Res}}({\mathcal{N}},{\mathfrak{C}})$ is a balanced network. Furthermore, there exists a multi-cycle flow ${\mathfrak{C}}$ such that the residual network has $c(e)=0$ for every $e\in E$.
@@
To prove the second part of the claim, consider some multi-cycle flow ${\mathfrak{C}}^{\prime}$ in ${\mathcal{N}}$. It is obvious that the capacity of every edge in ${\mathcal{N}}_{1}\coloneqq{\text{Res}}({\mathcal{N}},{\mathfrak{C}}^{\prime})$ is less than or equal to that in ${\mathcal{N}}$, with the capacity of at least one edge in ${\mathcal{N}}_{1}$ being strictly less than that in ${\mathcal{N}}$. If $c(e)>0$ for some $e\in E$ in ${\mathcal{N}}_{1}$, we can construct a simple cycle ${\mathcal{C}}_{1}$ containing $e$ by first constructing an edge cycle $E_{{\mathcal{C}}_{1}}$ explicitly starting from $e$; this is always possible because ${\mathcal{N}}_{1}$ is balanced. The magnitude of ${\mathcal{C}}_{1}$ is then given to be $\text{min}_{e\in E_{{\mathcal{C}}_{1}}}c(e)$. Notice this means the residual network ${\mathcal{N}}_{2}\coloneqq{\text{Res}}({\mathcal{N}}_{1},{\mathcal{C}}_{1})$ has at least one fewer edge than ${\mathcal{N}}_{1}$. We can then repeat this process with ${\mathcal{N}}_{1}$ and ${\mathcal{N}}$ and ${\mathcal{N}}_{2}$ in place of ${\mathcal{N}}_{1}$ and obtain a simple cycle flow ${\mathcal{C}}_{2}$. Repeating this until the resulting residual network has no edges, i.e. $c(e)=0$ for all edges, it then follows ${\mathfrak{C}}\coloneqq{\mathfrak{C}}^{\prime}\cup{\mathcal{C}}_{1}\cup\cdots\cup{\mathcal{C}}_{n}$ is a multi-cycle flow where the residual network ${\text{Res}}({\mathcal{N}},{\mathfrak{C}})$ has $c(e)=0$ for every edge. ∎
We now want to specialize even further and study a particular subclass of balanced networks that satisfies a crucial property known as nesting, which is the subclass of balanced networks where for any two disjoint subset of boundary vertices ${\underline{I}},{\underline{J}}$, there exists a multi-cycle flow ${\mathfrak{C}}({\underline{I}},{\underline{J}})$ that simultaneously maximizes the flux through both ${\underline{I}}$ and ${\underline{I}}{\underline{J}}$. Not every balanced network obeys the nesting condition, and we will prove a necessary and sufficient condition for nesting to hold. [^10] This is markedly different from the universal existence of nested flows for conventional flows involving sources and sinks (e.g., see [11]). It would be interesting to further explore connections between cycle flows and conventional flows, especially in the context of multi-commodity flows. As we will see in the next section, the nesting property will be repeatedly used to prove that (2.18) obeys both SA and SSA, thereby motivating us to identify it with the entropy.
@@
Given a balanced network ${\mathcal{N}}=(V,E,c)$ with boundary vertices $\partial V\subseteq V$, and given two disjoint subsets of boundary vertices ${\underline{I}},{\underline{J}}$ where ${\underline{I}}\cap{\underline{J}}=\emptyset$, let ${\underline{K}}\coloneqq\partial V\setminus{\underline{I}}{\underline{J}}$ denote the rest of the boundary vertices. Then there exists a multi-cycle flow ${\mathfrak{C}}({\underline{I}},{\underline{J}})$ with respect to the boundary vertex partition $\partial V=\{{\underline{I}},{\underline{J}},{\underline{K}}\}$, called a nested multi-cycle flow, whose magnitude through ${\underline{I}}$ and ${\underline{I}}{\underline{J}}$ is simultaneously maximized if and only if there exists a multi-cycle flow ${\mathfrak{C}}({\underline{I}}{\underline{J}})$ maximizing the flux through ${\underline{I}}{\underline{J}}$, a multi-cycle flow ${\mathfrak{C}}({\underline{I}})$ maximizing the flux through ${\underline{I}}$, and a multi-cycle flow ${\mathfrak{C}}$ that is a multi-cycle subflow of both ${\mathfrak{C}}({\underline{I}}{\underline{J}})$ and ${\mathfrak{C}}({\underline{I}})$ such that ${\underline{I}}$ and ${\underline{K}}$ are disjoint in ${\text{Res}}({\mathcal{N}},{\mathfrak{C}})$.
@@
First, let us prove the forward direction, and assume there exists a nested multi-cycle flow ${\mathfrak{C}}({\underline{I}},{\underline{J}})$ that maximizes the flux through both ${\underline{I}}$ and ${\underline{I}}{\underline{J}}$ simultaneously. We can then choose ${\mathfrak{C}}({\underline{I}})={\mathfrak{C}}({\underline{I}}{\underline{J}})={\mathfrak{C}}({\underline{I}},{\underline{J}})$. Then ${\mathfrak{C}}({\underline{I}},{\underline{J}})$ is trivially a multi-cycle subflow of both ${\mathfrak{C}}({\underline{I}})$ and ${\mathfrak{C}}({\underline{I}}{\underline{J}})$, so we can also let ${\mathfrak{C}}={\mathfrak{C}}({\underline{I}},{\underline{J}})$. We now claim that ${\underline{I}}$ and ${\underline{K}}$ are disjoint in the residual network ${\text{Res}}({\mathcal{N}},{\mathfrak{C}})$. To prove this claim, suppose the contrary and assume there is still a path between ${\underline{I}}$ and ${\underline{K}}$. Since ${\mathcal{N}}_{2}$ is balanced by 2, there still exists a simple cycle ${\mathcal{C}}$ between ${\underline{I}}$ and ${\underline{K}}$ (possibly including ${\underline{J}}$) in ${\text{Res}}({\mathcal{N}},{\mathfrak{C}})$. This means we can augment ${\mathfrak{C}}({\underline{I}},{\underline{J}})$ by ${\mathcal{C}}$ to increase the flux through ${\underline{I}}$, contradicting the fact ${\mathfrak{C}}({\underline{I}},{\underline{J}})$ maximizes the flux through ${\underline{I}}$. This proves ${\underline{I}}$ and ${\underline{K}}$ must be disjoint, thereby proving the forward direction.
@@
Next, we prove the reverse direction, and suppose there exists such a multi-cycle subflow ${\mathfrak{C}}$. In particular, this means in the residual network ${\text{Res}}({\mathcal{N}},{\mathfrak{C}})$, any vertex in ${\underline{J}}$ that is part of a cycle with a vertex in ${\underline{K}}$ cannot be connected to any vertices in ${\underline{I}}$, and similarly any vertex in ${\underline{J}}$ that is part of a cycle involving vertices in ${\underline{I}}$ cannot be connected to any vertex in ${\underline{K}}$. It follows that we can simultaneously construct multi-cycle flows ${\mathfrak{C}}_{{\underline{I}}{\underline{J}}}$ and ${\mathfrak{C}}_{{\underline{J}}{\underline{K}}}$ on ${\text{Res}}({\mathcal{N}},{\mathfrak{C}})$ such that ${\mathfrak{C}}_{{\underline{I}}{\underline{J}}}$ only involves cycles that maximize the flux between ${\underline{I}}$ and ${\underline{J}}$, and ${\mathfrak{C}}_{{\underline{J}}{\underline{K}}}$ only involves cycles that maximize the flux between ${\underline{J}}$ and ${\underline{K}}$. Thus, ${\mathfrak{C}}\cup{\mathfrak{C}}_{{\underline{I}}{\underline{J}}}\cup{\mathfrak{C}}_{{\underline{J}}{\underline{K}}}$ is a valid multi-cycle flow in the original network ${\mathcal{N}}$, and it utilizes all the capacities of edges between any two subsets ${\underline{I}},{\underline{J}}$, and ${\underline{K}}$. [^11] There might be additional cycle flows within a single subset, say ${\underline{I}}$, but such cycle flows do not contribute to the flux through ${\underline{I}},{\underline{J}}$, or ${\underline{K}}$ and hence can be ignored. It should then be clear that the flux through ${\mathfrak{C}}({\underline{I}}{\underline{J}})$ equals that through ${\mathfrak{C}}\cup{\mathfrak{C}}_{{\underline{J}}{\underline{K}}}$, as ${\mathfrak{C}}_{{\underline{I}}{\underline{J}}}$ does not contribute to the flux through ${\underline{I}}{\underline{J}}$. Similarly, the flux through ${\mathfrak{C}}({\underline{I}})$ must equal that through ${\mathfrak{C}}\cup{\mathfrak{C}}_{{\underline{I}}{\underline{J}}}$, as ${\mathfrak{C}}_{{\underline{J}}{\underline{K}}}$ does not contribute to the flux through ${\underline{I}}$. It follows that ${\mathfrak{C}}\cup{\mathfrak{C}}_{{\underline{J}}{\underline{K}}}\cup{\mathfrak{C}}_{{\underline{I}}{\underline{J}}}$ is precisely a multi-cycle flow that simultaneously maximizes the flux through both ${\underline{I}}$ and ${\underline{I}}{\underline{J}}$, completing the proof in the reverse direction as well. ∎
We conclude this section by restricting ourselves to studying the balanced networks that satisfy Theorem 1, which we define below.
Definition 9.
A balanced network ${\mathcal{N}}=(V,E,c)$ obeys nesting if for any pair of boundary vertex subsets ${\underline{I}},{\underline{J}}$, there exists a multi-cycle flow ${\mathfrak{C}}({\underline{I}},{\underline{J}})$ that simultaneously maximizes the flux through both ${\underline{I}}$ and ${\underline{I}}{\underline{J}}$. We will call such a network a nesting balanced network.
Note that not every balanced network is a nesting balanced network. An example of this is given in Figure 4. The maximal flux through $BC$ is 3, and the maximal flux through $ABCD$ is also 3. However, the multi-cycle flow depicted in Figure 4 is the only one that maximizes the flux through $ABCD$. However, it does not maximize the flux through $BC$, and so nesting fails for this balanced network. In fact, notice that this network violates SSA, as it is straightforward to check
${\texttt{S}}(ABC)+{\texttt{S}}(BCD)=3+2\ngeq{\texttt{S}}(BC)+{\texttt{S}}(ABCD)=3+3.$ (2.26)
In the next section, we will prove specifically for nesting balanced networks, SA and SSA are always satisfied.
$A$$B$$C$$D$$E$$F$